# LEKA: LLM-Enhanced Knowledge Augmentation

### v2

@@ 5.2 Experiment Results
The experiment results are shown in Table 3, where we compare our LEKA method across four datasets. In summary:
@@
(2) The LEKA outperforms all baseline methods in almost all metrics and datasets. Specifically, LEKA shows enhancements in accuracy by $2-5\%$ over other methods. It demonstrates notable improvements of up to $4.4\%$ in GBDT for the VID dataset and $5.4\%$ in RF for the BCW dataset. For improvement of overall accuracy, we demonstrate the LEKA’s effective retrieval and harmonization of feature space and marginal probability measures with the target datasets. LEKA demonstrates exceptional precision, improving the TT metric for the TCC dataset by over $10\%$ compared to baselines, effectively reducing false positives. Meanwhile, the LEKA consistently outperforms baselines in reducing misclassification rates with the highest F1 scores. These results prove that LEKA’s retrieval and data harmonization reduce the likelihood of misclassification by forming a deeper and clearer understanding of the potential relationship between features.
@@
We then compare our LEKA method with transfer learning methods to demonstrate its effectiveness in enhancing model performance in complex domains. In these scenarios, the transfer learning methods show their reliance on manual data selection and alignment processes. The results demonstrate that LEKA outperforms transfer learning methods FFT and TTab across various metrics and datasets. For example, in the BCW dataset, LEKA improves accuracy by $3.5\%$ and recall by $3.3\%$ compared to the competing methods. This superior performance is consistent across other datasets like VID, HD, and TCC, with notable enhancements in precision and recall, emphasizing LEKA’s effective data harmonization capabilities. This adaptability and efficiency in handling diverse datasets underscore LEKA’s robustness and advantages.
@@ 6 Conclusion
In this paper, we introduce LLM-Enhanced Knowledge Augmentation (LEKA), a novel retrieval and harmonization framework that dynamically refines source data for effective knowledge transfer. This structure significantly enhances the process of data augmentation by leveraging advanced LLM capabilities to automatically align and optimize data retrieved from diverse external libraries. Our extensive experiments across various tasks demonstrate the effectiveness and superiority over existing methods, especially in improving model adaptability and accuracy in complex data environments. By adopting LEKA on data-scarce domains, we achieve substantial improvements in learning performance and domain-specific task accuracy. For future work, we plan to extend this framework to include more varied data types and conduct a thorough empirical analysis to gain further insights into its underlying mechanisms and impacts.
@@ 7 Limitations
We acknowledge the following limitations of our work: (1) The current work has only been tested on tabular data, and more complex test scenarios have not yet been involved; (2) Despite its automation, LEKA can be computationally intensive. This could limit its applicability in resource-constrained environments or require substantial computational resources to maintain operational efficiency; (3) Adopting the LEKA in tasks with unique requirements may have inherent limitations.

### v3

@@ 5.2 Experiment Results
The results are shown in Table 3, comparing our LEKA across four datasets. In summary:
@@
(2) The LEKA outperforms all baseline methods in almost all metrics and datasets. Specifically, LEKA shows enhancements in accuracy by $2-5\%$ over other methods. It demonstrates notable improvements of up to $4.4\%$ in GBDT for the VID dataset and $5.4\%$ in RF for the BCW dataset. For improvement of overall accuracy, we demonstrate the LEKA’s effective retrieval and harmonization of feature space and marginal probability measures with the target datasets. LEKA demonstrates exceptional precision, improving the TT metric for the TCC dataset by over $10\%$ compared to baselines, effectively reducing false positives. Meanwhile, the LEKA consistently outperforms baselines in reducing misclassification rates with the highest F1 scores. These results prove that LEKA’s retrieval and data harmonization reduce misclassification by forming a deeper and clearer understanding of the potential relationship between features.
@@
We then compare our LEKA method with transfer learning methods to demonstrate its effectiveness in enhancing model performance in complex domains. In these scenarios, the transfer learning methods show their reliance on manual data selection and alignment processes. The results demonstrate that LEKA outperforms transfer learning methods, FFT, and TTab across various metrics and datasets. For example, in the BCW dataset, LEKA improves accuracy by $3.5\%$ and recall by $3.3\%$ compared to the competing methods. This superior performance is consistent across other datasets like VID, HD, and TCC, with notable enhancements in precision and recall, emphasizing LEKA’s effective data harmonization capabilities. This adaptability and efficiency in handling diverse datasets underscore LEKA’s robustness and advantages.
@@ 6 Conclusion
In this paper, we introduce LLM-Enhanced Knowledge Augmentation (LEKA), a novel retrieval and harmonization framework that dynamically refines source data for effective knowledge transfer. This structure significantly enhances data augmentation by leveraging advanced LLM capabilities to automatically align and optimize data retrieved from diverse external libraries. Extensive experiments across various tasks demonstrate superiority to existing methods, especially in improving model adaptability and accuracy in complex data environments. By adopting LEKA on data-scarce domains, we achieve substantial improvements in learning performance and domain-specific task accuracy. For future work, we plan to extend this framework to include more varied data types and conduct a thorough empirical analysis to understand its underlying mechanisms and impacts.
@@ 7 Limitations
We acknowledge the following limitations: (1) the current work has only been tested on tabular data, and more complex test scenarios have not yet been involved; (2) despite its automation, LEKA can be computationally intensive. This could limit its applicability in resource-constrained environments or require substantial computational resources to maintain operational efficiency; (3) adopting the LEKA in tasks with unique requirements may have inherent limitations.